# Learning in Markov Games with Adaptive Adversaries: Policy Regret, Fundamental Barriers, and Efficient Algorithms

### v1

@@ 5.2 Memory of any fixed length $m\geq 1$
6:     $\Pi^{k+1}:=\displaystyle\left\{\pi\in\Pi:\bar{V}^{\pi}(r_{{\mathcal{U}}^{k}},\hat{P}^{k},\Theta^{k})\geq\max_{\pi\in\Pi}\bar{V}^{\pi}(r_{{\mathcal{U}}^{k}},\hat{P}^{k},\Theta^{k})-cH^{2}SAB\sqrt{\alpha/(d^{*}T_{k})}\right\}$ where $r_{{\mathcal{U}}^{k}}(s_{1},a_{1},b_{1},\ldots,s_{H},a_{H},b_{H}):=\sum_{h\in[H]}1\{(h,s_{h},a_{h},b_{h},s_{h+1})\notin{\mathcal{U}}^{k}\}r_{h}(s_{h},a_{h},b_{h})$ and $\bar{V}^{\pi}(r,P,\Theta):={\textrm{OPTIMISTIC\_VALUE\_ESTIMATE}}(\pi,r,P,\Theta)$ is given in Algorithm 6
@@ A.1 Proof of Theorem 1
More formally, let’s consider any learner. Let ${\pi}^{1}$ be a policy that the learner commits in the first episode with the highest positive probability $p>0$. Note that $\pi^{1}$ and $p$ are the inherent property of the learner and do not depend on the adversary and the Markov game as in the first episode, the learner has zero information about the adversary and the Markov game. Now let’s consider the adversary that depends only on the learner’s policy in the first episode and nothing else, i.e., for all $t$ and policy sequence $\pi^{1},\ldots,\pi^{t}$, $f_{t}(\pi^{1},\ldots,\pi^{t})=f(\pi^{1})$ for some function $f:\Pi\rightarrow\Psi$. In addition, let $f$ such that $f(\pi)=\mu$ if $\pi=\pi^{1}$ and $f(\pi)=\nu$ otherwise, where $\mu$ and $\nu$ such that for all $s$, $\sup_{\pi}V_{1}^{\pi,\nu}(s)-\sup_{\pi}V_{1}^{\pi,\mu}(s)=\Omega(1)$. There exists a Markov game that always guarantees the existence of such $\mu,\nu$ (the constructions are fairly straightforward). Thus, with probability $p$, we have ${\textrm{PR}}(T)=\Omega(T)$. Note that the external regret $R(T)$ for this construction is 00.
@@ B.3.1 Sampling policies are sufficiently exploratory
$\sup_{\pi\in\Pi}\Pr[{\mathcal{F}}|P,\pi]\lesssim\frac{H^{3}\log(HSABK/\delta)}{T_{k}}+H\xi_{MLE}(T_{k}).$
@@ B.3.2 Uniform policy evaluation
Conditioned on the event $E$ in Lemma B.5 and the event $E_{k}$ in Lemma B.8 and the high-probability event in Lemma B.9, with probability at least $1-\delta$, for any $k\in[K]$, any reward function $r$, and any policy $\pi\in\Pi$, we have
@@
$\geq\sup_{\pi\in\Pi}\bar{V}^{\pi}(r_{{\mathcal{U}}^{k}},\hat{P}^{k},\Theta^{k})-{\mathcal{O}}\left(H^{2}SAB\sqrt{\frac{\alpha}{d^{*}T_{k}}}+HS^{3/2}AB\sqrt{\frac{\log(HAT/\delta)}{T_{k}}}\right)$
$\geq\sup_{\pi\in\Pi}V^{\pi,f([\pi]^{m})}(r_{{\mathcal{U}}^{k}},\hat{P}^{k})-{\mathcal{O}}\left(H^{2}SAB\sqrt{\frac{\alpha}{d^{*}T_{k}}}+HS^{3/2}AB\sqrt{\frac{\log(HAT/\delta)}{T_{k}}}\right)$
$\geq\sup_{\pi\in\Pi}V^{\pi,f([\pi]^{m})}(r_{{\mathcal{U}}^{k}},\tilde{P}^{k})-{\mathcal{O}}\left(H^{2}SAB\sqrt{\frac{\alpha}{d^{*}T_{k}}}+HS^{3/2}AB\sqrt{\frac{\log(HAT/\delta)}{T_{k}}}\right)$
@@
where the first inequality follows from the first part of Lemma B.10, the second inequality follows from Lemma B.12, the third inequality follows from Lemma B.11, the fourth inequality follows from the definition of $\Pi^{k+1}$, the fifth inequality follows from Lemma B.11, the sixth inequality follows from Lemma B.12, and the last inequality follows from the second part of Lemma B.10. ∎

### v2

@@ 5.2 Memory of any fixed length $m\geq 1$
6:     $\Pi^{k+1}:=\displaystyle\left\{\pi\in\Pi^{k}:\bar{V}^{\pi}(r_{{\mathcal{U}}^{k}},\hat{P}^{k},\Theta^{k})\geq\max_{\pi\in\Pi^{k}}\bar{V}^{\pi}(r_{{\mathcal{U}}^{k}},\hat{P}^{k},\Theta^{k})-cH^{2}SAB\sqrt{\alpha/(d^{*}T_{k})}\right\}$ where $r_{{\mathcal{U}}^{k}}(s_{1},a_{1},b_{1},\ldots,s_{H},a_{H},b_{H}):=\sum_{h\in[H]}1\{(h,s_{h},a_{h},b_{h},s_{h+1})\notin{\mathcal{U}}^{k}\}r_{h}(s_{h},a_{h},b_{h})$ and $\bar{V}^{\pi}(r,P,\Theta):={\textrm{OPTIMISTIC\_VALUE\_ESTIMATE}}(\pi,r,P,\Theta)$ is given in Algorithm 6
@@ A.1 Proof of Theorem 1
More formally, let’s consider any learner. Let ${\pi}^{1}$ be a policy that the learner commits in the first episode with the highest positive probability $p>0$. Note that $\pi^{1}$ and $p$ are the inherent property of the learner and do not depend on the adversary and the Markov game as in the first episode, the learner has zero information about the adversary and the Markov game. Now let’s consider the adversary that depends only on the learner’s policy in the first episode and nothing else, i.e., for all $t$ and policy sequence $\pi^{1},\ldots,\pi^{t}$, $f_{t}(\pi^{1},\ldots,\pi^{t})=f(\pi^{1})$ for some function $f:\Pi\rightarrow\Psi$. In addition, let $f$ such that $f(\pi)=\mu$ if $\pi=\pi^{1}$ and $f(\pi)=\nu$ otherwise, where $\mu$ and $\nu$ such that for all $s$, $\sup_{\pi\neq\pi^{1}}V_{1}^{\pi,\nu}(s)-\sup_{\pi}V_{1}^{\pi,\mu}(s)=\Omega(1)$. There exists a Markov game that always guarantees the existence of such $\mu,\nu$ (the constructions are fairly straightforward). Thus, with probability $p$, we have ${\textrm{PR}}(T)=\Omega(T)$. Note that the external regret $R(T)$ for this construction is 00.
@@ B.3.1 Sampling policies are sufficiently exploratory
$\sup_{\pi\in\Pi^{k}}\Pr[{\mathcal{F}}|P,\pi]\lesssim\frac{H^{3}\log(HSABK/\delta)}{T_{k}}+H\xi_{MLE}(T_{k}).$
@@ B.3.2 Uniform policy evaluation
Conditioned on the event $E$ in Lemma B.5 and the event $E_{k}$ in Lemma B.8 and the high-probability event in Lemma B.9, with probability at least $1-\delta$, for any $k\in[K]$, any reward function $r$, and any policy $\pi\in\Pi^{k}$, we have
@@
Lemma B.14.
Let $\pi^{*}=\operatorname*{arg\,max}_{\pi\in\Pi}V^{\pi,f([\pi]^{m})}$. Conditioned on the event of Lemma B.10, Lemma B.11, Lemma B.12, $\pi^{*}$ never get eliminated from $\Pi^{k}$ for $k\in[K]$ by Algorithm 3.
Proof of Lemma B.14.
We will prove by induction. Since $\Pi^{1}$ contains all the possible policies of the learner, $\pi^{*}\in\Pi^{1}$. Assume by induction that $\pi^{*}\in\Pi^{k}$. We will show that $\pi^{*}\in\Pi^{k+1}$. Indeed, let $\hat{\pi}^{k}=\operatorname*{arg\,max}_{\pi\in\Pi^{k}}\bar{V}^{\pi}(r_{{\mathcal{U}}^{k}},\hat{P}^{k},\Theta^{k})$, we have
[…]
$-\frac{H^{4}\log(HSABK/\delta)}{T_{k}}-H^{2}\xi_{MLE}(T_{k})$
$\overset{\pi^{*}\text{ is optimal}}{\geq}V^{\hat{\pi}^{k},f([\hat{\pi}^{k}]^{m})}(r,P)-HS^{3/2}AB\sqrt{\frac{\log(HAT/\delta)}{T_{k}}}-HSAB\cdot\xi_{MLE}(T_{k})$
$-\frac{H^{4}\log(HSABK/\delta)}{T_{k}}-H^{2}\xi_{MLE}(T_{k})$
[…]
$-\frac{H^{4}\log(HSABK/\delta)}{T_{k}}-H^{2}\xi_{MLE}(T_{k})$
[…]
$-\frac{H^{4}\log(HSABK/\delta)}{T_{k}}-H^{2}\xi_{MLE}(T_{k})$
[…]
$-\frac{H^{4}\log(HSABK/\delta)}{T_{k}}-H^{2}\xi_{MLE}(T_{k})-H^{2}\sqrt{\frac{\alpha}{d^{*}T_{k}}}$
$\geq\max_{\pi\in\Pi^{k}}V^{\pi,f([\pi]^{m})}(r_{{\mathcal{U}}^{k}},\hat{P}^{k},\Theta^{k})-{\mathcal{O}}\bigg{(}H^{2}(SAB+H)\sqrt{\frac{\alpha}{d^{*}T_{k}}}+\frac{H^{4}\log(HSABK/\delta)}{T_{k}}$
$+HS^{3/2}AB\sqrt{\frac{\log(HAT/\delta)}{T_{k}}}\bigg{)}.$
Thus, $\pi^{*}\in\Pi^{k+1}$. ∎
@@
$\geq\sup_{\pi\in\Pi^{k}}\bar{V}^{\pi}(r_{{\mathcal{U}}^{k}},\hat{P}^{k},\Theta^{k})-{\mathcal{O}}\left(H^{2}SAB\sqrt{\frac{\alpha}{d^{*}T_{k}}}+HS^{3/2}AB\sqrt{\frac{\log(HAT/\delta)}{T_{k}}}\right)$
$\geq\sup_{\pi\in\Pi^{k}}V^{\pi,f([\pi]^{m})}(r_{{\mathcal{U}}^{k}},\hat{P}^{k})-{\mathcal{O}}\left(H^{2}SAB\sqrt{\frac{\alpha}{d^{*}T_{k}}}+HS^{3/2}AB\sqrt{\frac{\log(HAT/\delta)}{T_{k}}}\right)$
$\geq\sup_{\pi\in\Pi^{k}}V^{\pi,f([\pi]^{m})}(r_{{\mathcal{U}}^{k}},\tilde{P}^{k})-{\mathcal{O}}\left(H^{2}SAB\sqrt{\frac{\alpha}{d^{*}T_{k}}}+HS^{3/2}AB\sqrt{\frac{\log(HAT/\delta)}{T_{k}}}\right)$
$\geq\sup_{\pi\in\Pi^{k}}V^{\pi,f([\pi]^{m})}(r,P)$
@@
where the first inequality follows from the first part of Lemma B.10, the second inequality follows from Lemma B.12, the third inequality follows from Lemma B.11, the fourth inequality follows from the definition of $\Pi^{k+1}$, the fifth inequality follows from Lemma B.11, the sixth inequality follows from Lemma B.12, and the seventh inequality follows from the second part of Lemma B.10, and the last inequality follows from Lemma B.14. ∎